# Space Group Informed Transformer for Crystalline Materials Generation

### v1

@@ IV.2 Related works
Crystal generative models have been explored using variational autoencoder [72, 23], generative adversarial networks [73, 12], normalizing flows [74, 75, 76], diffusion models [11, 12, 13, 26, 14, 15, 16, 21, 22], GFlowNet [24, 25], and autoregressive models [77, 55, 78, 17, 18, 19, 19]. In these autoregressive models, one either use atomistic features [77, 55, 78] or use pure text tokens [17, 18, 19, 20]. However, with the introduction of specialized tokens for crystals, the boundary between the two is blurred.
@@
The CrystalFormer is most closely related to the autoregressvie generative model originally designed for molecules [77, 55, 78]. However, instead of predicting the relative distances of atoms, we predict the Wyckoff positions of atoms in the unit cell. Having the luxury of the space group symmetry for crystals provides strong hints on where to put the atoms in the unit cell and greatly simplifies the design around rotational equivariance. On the other hand, compared to Refs. [17, 18, 19, 20] which treat text descriptions of crystals using autoregressive language models, we are only dealing with more concise and essential atomistic representation of crystals, which leads to smaller model size and faster sampling speed. Fast generation speed is not only a welcoming feature but also will be crucial for intelligent exploration of materials space based on combinations of probabilistic generation and post-selection, backtracking, and searching techniques [79]. More importantly, CrystalFormer guarantees space group constraints since these symmetries are baked in the model architecture rather than learned as statistical correlation from text data. In a sense, the present work employs intrinsic mathematical (as opposed to natural) language to incorporate the symmetry principle in the generative modeling of crystals.
@@
As a side remark, the Wyckoff position features have been used in machine learning models for materials property prediction [80, 81]. Incorporating space group information in the encoder-only transformer models may also boost their property prediction performance [82, 83] as suggested by the related study [84].
V Outlook
@@
Crystal structure prediction has long been the dream of computational material sciences [85]. CrystalFormer integrates exact symmetry principles from math and empirical chemical intuitions from data into one framework. Probabilistic generative modeling of crystalline materials using CrystalFormer opens the way to many future innovations in materials discovery.
@@
Precisely controlling the space group in the generative model of crystalline materials is not only a highly desired feature but also greatly simplifies the task. An obvious future direction is to scale up the model as well as the training dataset, especially curating a dataset with better coverage of space groups. Note that the MP-20 dataset has by no means exhausted all available crystalline material [19, 16]. The transformer-based generative model is ready to be scaled up with more training data, in the same fashion as large language models [86]. Given similar model architectures, the idea of generative pretraining of a foundational model for material generation is appealing. When scaling up the model it will be interesting to note the possible appearance of neural scaling law [87] as it has also been showing up in other contexts of atomistic modeling [88].
@@
The model architecture is still open to further improvement to better serve the purpose of material discovery, First of all, to better facilitate data efficiency learning and structure phase transitions-related applications, it will be useful to further exploit the group-subgroup structure [89] in the model architecture. Second, besides chemical elements, it will be useful to constrain the stoichiometry of generated materials. This may be achieved by tuning the logit bias or combining the present probabilistic generation approaches with traditional enumeration approaches [90]. Lastly, it is worth exploring to use CrystalFormer as the base distribution in the flow model and employ symmetry-persevering transportation’s to further adjust the atoms coordinates and unit cells [10, 26], which mimics a symmetry-constrained relaxation process [91].
Conditioned crystal structure generation based on material properties [78, 16] and experimental measurements [92] are both straightforward with CrystalFormer. One can either extend the space group embedding or employ the full encoder-decoder transformer architecture [36] for conditioned crystal generation. The first step of property-guided generation could take into account of synthesizability and stability of materials. In the future, proposing materials with multiple properties is a promising direction for material generative models. To achieve that goal, curating a high-quality material dataset with either computed or measured properties is crucial. Needless to say, to close the loop of material discovery, one will need to carry out experimental verification for the generated materials.
@@ Acknowledgements.
We thank Han Wang, Lin Yao, Linfeng Zhang, Chen Fang, Yanchao Wang, Qi Yang, Shigang Ou, Xinyang Dong, and Quansheng Wu for useful discussions. This project is supported by the National Natural Science Foundation of China under Grants No. T2225018, No. 92270107, No. 12188101, No. T2121001, and No. 12034009 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000 and No. XDB30000000.References
@@
- Ye et al. [2024] C.-Y. Ye, H.-M. Weng, and Q.-S. Wu, Con-cdvae: A method for the conditional generation of crystal structures, (2024), arXiv:2403.12478 [cond-mat.mtrl-sci] .
- Zhu et al. [2023] R. Zhu, W. Nong, S. Yamazaki, and K. Hippalgaonkar, Wycryst: Wyckoff inorganic crystal generator framework, (2023), arXiv:2311.17916 [cond-mat.mtrl-sci] .
@@
- Note [1] We follow the convention that capital letters appear after lower case letters. This handles the edge case of the No. 47 space group whose Wyckoff positions reach ’A’. This rule also demands that the padded atoms appear at the end of the sequence with Wyckoff letter ’X’.
@@
- Wang et al. [2021a] H.-C. Wang, S. Botti, and M. A. L. Marques, Predicting stable crystalline compounds using chemical similarity, npj Computational Materials 7, 12 (2021a).
@@
- [56] https://github.com/sparks-baird/matbench-genmetrics.
- Note [2] Due to merging of Wyckoff positions that are occupied with the same elements, 26 structures actually belong to the space group (No. 221).
- Jain et al. [2013] A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, and K. A. Persson, Commentary: The Materials Project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
- Wang et al. [2012] Y. Wang, J. Lv, L. Zhu, and Y. Ma, Calypso: A method for crystal structure prediction, Computer Physics Communications 183, 2063 (2012).
@@
- Wang et al. [2021b] H.-C. Wang, S. Botti, and M. A. L. Marques, Predicting stable crystalline compounds using chemical similarity, npj Computational Materials 7, 12 (2021b).
- Schmidt et al. [2022a] J. Schmidt, N. Hoffmann, H.-C. Wang, P. Borlido, P. J. M. A. Carriço, T. F. T. Cerqueira, S. Botti, and M. A. L. Marques, Large-scale machine-learning-assisted exploration of the whole materials space (2022a), arXiv:2210.00579 [cond-mat.mtrl-sci] .
- Schmidt et al. [2022b] J. Schmidt, H.-C. Wang, T. F. T. Cerqueira, S. Botti, and M. A. L. Marques, A dataset of 175k stable and metastable materials calculated with the pbesol and scan functionals, Scientific Data 9, 64 (2022b).
@@ Appendix A More details of CrystalFormer
To recap, the space group information plays a key role in the architecture, training, and sampling of CrystalFormer. First of all, the vector embedding of space group number controls all subsequent outputs in the transformer that determines the Wyckoff letters, chemical species, fractional coordinates, and lattice parameters. Second, the information of the space group and Wyckoff letter are used to select active components in the fractional coordinates and lattice parameters in the loss function during training. Lastly, the space group determines the concrete meaning of in terms of multiplicities and fractional coordinates. This information is used to place right amount of atoms to appropriate locations during sampling.
@@ A.2 Sampling algorithm
Algorithm 2 summarizes the sampling method of CrystalFormer. It takes 520 seconds to generate a batch size 13,000 crystal samples on a single A100 GPU, which translates to a generation speed 40 millisecond per sample.
Appendix B Stable crystal samples
Table S2 list generated samples in the space group (No. 225) with . Three of them (InNd, InTb, and EuTl) actually fall into the space group (No. 221) due to merging of ”4a” and ”4b” Wyckoff positions that are occupied by the same elements. Overall, out of 1000 samples generated by CrystalFormer, we found 10 stable materials that are not contained in the Materials Project database. Extending the search to more recent material databases, we found match of AcCuRh, AcHgIn, AcAgSi and AuInPm in the WBM dataset [93], AcHgIn, AcAgSi, InPmTl, PdPmZn, IrLuPm and AuInPm in the Alexandria dataset [94, 95]. Neverthless, we did not find AcBiK, InTb, and CaPaRu to be present in any of these dataset.

### v2

@@ V Related works
Crystal generative models have been explored using variational autoencoder [93, 23], generative adversarial networks [94, 12], normalizing flows [95, 96, 97, 98], diffusion models [11, 12, 13, 26, 14, 15, 16, 21, 22], GFlowNet [24, 25], and autoregressive models [99, 100, 101, 17, 18, 19, 20]. In these autoregressive models, one either uses atomistic features [99, 100, 101, 17] or uses pure text descriptions [18, 19, 20]. Nevertheless, with the introduction of specialized tokens for crystals, the boundary between the two is blurred.
@@
The CrystalFormer is most closely related to the autoregressive generative model originally designed for molecules [99, 100, 101]. However, instead of predicting the relative distances of atoms, we predict the Wyckoff positions of symmetry-inequivalent atoms in the unit cell. Having the luxury of the space group symmetry for crystals provides strong hints on where to put the atoms in the unit cell and greatly simplifies the design around spatial symmetries. On the other hand, compared to Refs. [19] which treats text descriptions of crystals using autoregressive language models, we are only dealing with a more concise and essential atomistic representation of crystals, which leads to a smaller model size and faster sampling speed. Fast generation speed is not only a welcoming feature but also will be crucial for further exploration of materials space based on combinations of probabilistic generation and post-selection, Monte Carlo sampling, backtracking, and searching techniques [102]. More importantly, by baking in the space group symmetry in the model rather than learning them as statistical correlation from texts [18, 20], CrystalFormer guarantees space group constraints and cherishes the precious data and computing time. In a sense, the present work employs intrinsic mathematical (as opposed to natural) language to incorporate the symmetry principle in the generative modeling of crystals.
@@
As a side remark, the Wyckoff position features have been used in machine learning models for materials property prediction [49, 103]. Incorporating space group information in the encoder-only transformer models may also enhance their property prediction performance [104, 105, 106] as suggested by Ref. [107].
VI Outlook
@@
Precisely controlling the space group in the generative model of crystalline materials not only greatly simplifies the task but also is a highly desired feature for materials discovery and design. CrystalFormer integrates exact symmetry principles from math and empirical chemical intuitions from data into one unified framework. Probabilistic generative modeling of crystalline materials using CrystalFormer opens the way to many future innovations in materials discovery.
@@
Note that the MP-20 dataset has by no means exhausted all available crystalline material [19, 16]. An obvious future direction is to scale up the model as well as the training dataset, especially curating a dataset with better coverage of space groups. In particular, extending the dataset to include both inorganic and organic crystals [108] may be beneficial as it improves the data coverage of low symmetric space groups. The transformer-based generative model is ready to be scaled up to work with much larger and more diverse training data, in the same fashion as large language models [109]. Given similar model architectures, the idea of generative pretraining of a foundational model for material generation is appealing. When scaling up the model it will be interesting to note the possible appearance of neural scaling law [110] as it has also been showing up in other contexts of atomistic modeling [111].
@@
The model architecture and sampling strategy are both open to further refinement to better serve the purpose of material discovery. First of all, to better facilitate data efficiency learning and structure phase transitions-related applications, it will be useful to further exploit the Euclidean normalizer [112] and group-subgroup relation [113] in the model architecture or training procedure. Second, it is worth exploring using CrystalFormer as the base distribution in the flow model and employs symmetry-persevering transportation to further adjust the atoms coordinates and unit cells [10, 26], which mimics a symmetry-constrained relaxation process [114]. Lastly, it may be worth employing more advanced constrained and guided sequence generation methods [115, 116, 117, 118] for more flexible control on the elements, structure, or stoichiometry of generated materials.
Conditioned materials generation depending on properties [101, 16, 21, 22] and experimental measurements [119] are highly desired features of materials generative model. Although it is straightforward to extend CrystalFormer (e.g. extend the space group embedding or employ the encoder-decoder transformer architecture [40]) to incorporate these conditions, we are particularly excited about the plug-and-play routine demonstrated in Sec. IV.3. Along this line, we envision an ecosystem [120] where the foundational generative model for and more specialized discriminative models are developed separately but brought together via the Bayes rule.
@@ Acknowledgements.
We thank Han Wang, Lin Yao, Linfeng Zhang, Chen Fang, Yanchao Wang, Zhenyu Wang, Qi Yang, Shigang Ou, Xinyang Dong, Wenbing Huang, Quansheng Wu, Wanjian Yin, Xi Dai, Shuang Jia, Hangtian Zhu, Jiangang Guo, and Hongjian Zhao for useful discussions. This project is supported by the National Natural Science Foundation of China under Grants No. T2225018, No. 92270107, No. 12188101, No. T2121001, and No. 12034009 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000 and No. XDB30000000.References
@@
- Ye et al. [2024a] C.-Y. Ye, H.-M. Weng, and Q.-S. Wu, Con-cdvae: A method for the conditional generation of crystal structures, (2024a), arXiv:2403.12478 [cond-mat.mtrl-sci] .
- Zhu et al. [2024] R. Zhu, W. Nong, S. Yamazaki, and K. Hippalgaonkar, Wycryst: Wyckoff inorganic crystal generator framework, Matter doi.org/10.1016/j.matt.2024.05.042 (2024).
@@
- Cheetham and Seshadri [2024] A. K. Cheetham and R. Seshadri, Artificial intelligence driving materials discovery? perspective on the article: Scaling deep learning for materials discovery, Chemistry of Materials 36, 3490 (2024).
- Hornfeck [2022] W. Hornfeck, On the combinatorics of crystal structures: number of wyckoff sequences of given length, Acta Crystallographica Section A: Foundations and Advances 78, 149 (2022).
@@
- Liang et al. [2020] H. Liang, V. Stanev, A. G. Kusne, and I. Takeuchi, Cryspnet: Crystal structure predictions via neural networks, Phys. Rev. Mater. 4, 123802 (2020).
@@
- Venkatraman and Carvalho [2024] V. Venkatraman and P. A. Carvalho, Accurate space-group prediction from composition, Journal of Applied Crystallography (2024).
@@
- Note [1] We follow the convention that capital letters appear after lower case letters. This handles the edge case of the space group (No. 47) whose Wyckoff letters used up 26 lowercase letters and reached ’A’ for the generic position. In addition, we use the letter ’X’ to indicate the Wyckoff position of padding atoms that appear at the end of the sequence, see e.g. of Fig. 1(b).
@@
- Hautier et al. [2011] G. Hautier, C. Fischer, V. Ehrlacher, A. Jain, and G. Ceder, Data mined ionic substitutions for the discovery of new compounds, Inorganic chemistry 50, 656 (2011).
@@
- Wang et al. [2021] H.-C. Wang, S. Botti, and M. A. L. Marques, Predicting stable crystalline compounds using chemical similarity, npj Computational Materials 7, 12 (2021).
@@
- Gazzarrini et al. [2024] E. Gazzarrini, R. K. Cersonsky, M. Bercx, C. S. Adorf, and N. Marzari, The rule of four: anomalous distributions in the stoichiometries of inorganic compounds, Npj Computational Materials 10, 73 (2024).
- Palgrave [2024] R. Palgrave, An explanation for the rule of four in inorganic materials 10.26434/chemrxiv-2024-sxqwh (2024).
- Alvarez [2005] S. Alvarez, Polyhedra in (inorganic) chemistry, Dalton Transactions , 2209 (2005).
- Regnault et al. [2022] N. Regnault, Y. Xu, M.-R. Li, D.-S. Ma, M. Jovanovic, A. Yazdani, S. S. Parkin, C. Felser, L. M. Schoop, N. P. Ong, et al., Catalogue of flat-band stoichiometric materials, Nature 603, 824 (2022).
@@
- Wales [1998] D. J. Wales, Symmetry, near-symmetry and energetics, Chemical physics letters 285, 330 (1998).
@@
- Wang et al. [2010] Y. Wang, J. Lv, L. Zhu, and Y. Ma, Crystal structure prediction via particle-swarm optimization, Phys. Rev. B 82, 094116 (2010).
- Falls et al. [2020] Z. Falls, P. Avery, X. Wang, K. P. Hilleke, and E. Zurek, The xtalopt evolutionary algorithm for crystal structure prediction, The Journal of Physical Chemistry C 125, 1601 (2020).
@@
- Zhang et al. [2024] Q. Zhang, A. Choudhury, and A. Chernatynskiy, A symmetry-oriented crystal structure prediction method for crystals with rigid bodies, (2024), arXiv:2407.21337 [cond-mat.mtrl-sci] .
@@
- Vasala and Karppinen [2015] S. Vasala and M. Karppinen, A2b’b”o6 perovskites: a review, Progress in solid state chemistry 43, 1 (2015).
- Wang et al. [2024] Y. Wang, B. Baldassarri, J. Shen, J. He, and C. Wolverton, Landscape of thermodynamic stabilities of a2bb’o6 compounds, Chemistry of Materials (2024).
- Miao et al. [2019] N. Miao, H. Zhou, L. Mou, R. Yan, and L. Li, CGMH: Constrained sentence generation by metropolis-hastings sampling, in Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 33 (2019) pp. 6834–6842.
- Chu et al. [2018] I.-H. Chu, S. Roychowdhury, D. Han, A. Jain, and S. P. Ong, Predicting the volumes of crystals, Computational Materials Science 146, 184 (2018).
- Okabe et al. [2024] R. Okabe, M. Cheng, A. Chotrattanapituk, N. T. Hung, X. Fu, B. Han, Y. Wang, W. Xie, R. J. Cava, T. S. Jaakkola, Y. Cheng, and M. Li, Structural constraint integration in generative model for discovery of quantum material candidates, (2024), arXiv:2407.04557 [cond-mat.mtrl-sci] .
- Verkuil et al. [2022] R. Verkuil, O. Kabeli, Y. Du, B. I. Wicky, L. F. Milles, J. Dauparas, D. Baker, S. Ovchinnikov, T. Sercu, and A. Rives, Language models generalize beyond natural proteins, BioRxiv , 2022 (2022).
- Meredig and Wolverton [2013] B. Meredig and C. Wolverton, A hybrid computational–experimental approach for automated crystal structure solution, Nature materials 12, 123 (2013).
- Parackal et al. [2023] A. S. Parackal, R. E. A. Goodall, F. A. Faber, and R. Armiento, Identifying crystal structures beyond known prototypes from x-ray powder diffraction spectra, (2023), arXiv:2309.16454 [cond-mat.mtrl-sci] .
- Xie and Grossman [2018] T. Xie and J. C. Grossman, Crystal graph convolutional neural networks for an accurate and interpretable prediction of material properties, Phys. Rev. Lett. 120, 145301 (2018).
- Ye et al. [2024b] C.-Y. Ye, H.-M. Weng, and Q.-S. Wu, Con-cdvae: A method for the conditional generation of crystal structures, Computational Materials Today 1, 100003 (2024b).
- Franceschetti and Zunger [1999] A. Franceschetti and A. Zunger, The inverse band-structure problem of finding an atomic configuration with given electronic properties, Nature 402, 60 (1999).
- Cheng et al. [2023] G. Cheng, X.-G. Gong, and W.-J. Yin, Global optimization in the discrete and variable-dimension conformational space: The case of crystal with the strongest atomic cohesion, (2023), arXiv:2302.13537 [cond-mat.mtrl-sci] .
- Note [2] For simplicity of the Wyckoff sequence are kept unchanged in the MCMC sampling.
- Zhao et al. [2017] X.-G. Zhao, J.-H. Yang, Y. Fu, D. Yang, Q. Xu, L. Yu, S.-H. Wei, and L. Zhang, Design of lead-free inorganic halide perovskites for solar cells via cation-transmutation, Journal of the American Chemical Society 139, 2630 (2017).
@@
- Miller et al. [2024] B. K. Miller, R. T. Chen, A. Sriram, and B. M. Wood, Flowmm: Generating materials with riemannian flow matching, (2024), arXiv:2406.04713 [cs.CL] .
@@
- Xu et al. [2024] H. Xu, D. Qian, and J. Wang, Predicting many properties of crystals by a single deep learning model, (2024), arXiv:2405.18944 [cond-mat.mtrl-sci] .
@@
- Groom et al. [2016] C. R. Groom, I. J. Bruno, M. P. Lightfoot, and S. C. Ward, The cambridge structural database, Structural Science 72, 171 (2016).
@@
- Dathathri et al. [2019] S. Dathathri, A. Madotto, J. Lan, J. Hung, E. Frank, P. Molino, J. Yosinski, and R. Liu, Plug and play language models: A simple approach to controlled text generation, (2019), arXiv:1912.02164 [cs.CL] .
- Zhang et al. [2020] M. Zhang, N. Jiang, L. Li, and Y. Xue, Language generation via combinatorial constraint satisfaction: A tree search enhanced monte-carlo approach, (2020), arXiv:2011.12334 [cs.LG] .
- Qin et al. [2022] L. Qin, S. Welleck, D. Khashabi, and Y. Choi, Cold decoding: Energy-based constrained text generation with langevin dynamics, Advances in Neural Information Processing Systems 35, 9538 (2022).
- Lew et al. [2023] A. K. Lew, T. Zhi-Xuan, G. Grand, and V. K. Mansinghka, Sequential monte carlo steering of large language models using probabilistic programs, (2023), arXiv:2306.03081 [cs.AI] .
@@
- Raymond [1999] E. S. Raymond, The Cathedral and the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary (O’Reilly Media, Sebastopol, CA, 1999).
- [123] https://github.com/sparks-baird/matbench-genmetrics.
@@
- Chen et al. [2020] M. Chen, A. Radford, R. Child, J. Wu, H. Jun, D. Luan, and I. Sutskever, Generative pretraining from pixels, in Proceedings of the 37th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol. 119, edited by H. D. III and A. Singh (PMLR, 2020) pp. 1691–1703.
@@ Appendix A More details of CrystalFormer
To recap, the space group information plays a key role in the architecture, training, and sampling of CrystalFormer. First of all, the vector embedding of space group number controls all subsequent outputs of the transformer corresponding to the Wyckoff letters, chemical species, fractional coordinates, and lattice parameters. Second, the information of the space group and Wyckoff letter are used to select active components in the fractional coordinates and lattice parameters in the loss function during training. Lastly, the space group determines the concrete meaning of Wyckoff letters in terms of multiplicities and fractional coordinates, which are used to place the right number of atoms precisely in the unit cell during sampling.
@@ A.2 Sampling algorithm
Algorithm 2 summarizes the autoregressive sampling method of CrystalFormer. It takes 520 seconds to generate a batch size 13,000 crystal samples on a single A100 GPU, which translates to a generation speed of 40 milliseconds per sample.
Algorithm 3 summarizes the Markov chain Monte Carlo sampling of CrystalFormer. It is used in the structure constrained generation of crystals in Sec. IV.2. The property-guided materials design discussed in Sec. IV.3 employs a similar sampling strategy for the posterior probability distribution Eq. (3).
Appendix B Validity and novelty of generated samples
Figure S2 illustrates the structure and compositional validity of generated samples across all 230 space groups. Following the Ref. [121], a structure meets the validity criteria if the shortest atomic distance exceeds , a lenient standard. Composition validity requires charge neutrality as computed by SMACT [122]. This is, however, an overly stringent criterion since the composition validity of the training set is only around by this measure [11]. Note that the CrystalFormer can generate reasonable samples even for those space groups without any training data. This because the model can exploit knowledge learned from other space groups to place suitable atoms in the Wyckoff positions due to weight sharing. Moreover, since the sampling process makes use the of Wyckoff position table. The three dimensional coordinates of atoms are not completely random even for unseen space groups.
| Space group | Crystal system | Training samples | Validity () | |
| Struc. | Comp. | |||
| 2 | Triclinic | 676 | 83.10 | 83.0 |
| 12 | Monoclinic | 1273 | 87.70 | 81.80 |
| 62 | Orthorhombic | 1187 | 95.50 | 87.20 |
| 139 | Tetragonal | 1233 | 97.70 | 83.40 |
| 166 | Trigonal | 1076 | 98.50 | 85.0 |
| 194 | Hexagonal | 1129 | 99.40 | 89.90 |
| 225 | Cubic | 3960 | 99.60 | 93.50 |
| 1 | Triclinic | 27136 | 91.40 | 80.20 |
| Autoregressive models | ||||
| PGSchNet [100] | 99.65 | 75.96 | ||
| LM-CH (character-level tokenization) [18] | 84.81 | 83.55 | ||
| LM-AC (atom coordinate-level tokenization) [18] | 95.81 | 88.87 | ||
| Crystal-text-LLM [20] | 96.5 | 86.3 | ||
| Diffusion models | ||||
| CDVAE [11] | 100.0 | 86.70 | ||
| DiffCSP [13] | 100.0 | 83.25 | ||
| DiffCSP++ [26] | 99.94 | 85.12 | ||
| UniMat-Large [15] | 97.2 | 89.4 | ||
Table LABEL:table:validity_table reports the validity of generated samples for selected space groups in each of the seven crystal systems. To ensure the numbers are representative, we chose the space group to be the one with the most training data for each crystal system. One sees that the model performs better for more symmetric space groups. This is a nice feature that complements existing crystal generative models, which mostly have difficulties in generating highly symmetric structures. As a reference, we also list the validity of the generated samples suppose one treats the crystals as if they are all in the space group (No. 1) with only translational symmetry. One sees the structure validity scores of space group improve compared to the one shown in Figure S2 due to increased training samples.
The second part of Table LABEL:table:validity_table shows reference results in the literature for the same validity test. In principle, the performance of the present model should fall back to the language model approaches [18, 20]. The remaining gap may be due to details such as including the header line in the crystallographic information files (CIF), specific sampling strategy of language models, or the additional post-selection of samples [20]. In the table, the DiffCSP++ [26] is the only alternative model that exploits the space group symmetry in a rigorous manner similar to ours. However, the DiffCSP++ does not predict Wyckoff positions in the generation process. Instead, one needs to search for template structures in the training set for generation, which may limit its generality. Besides works listed in Table LABEL:table:validity_table that reported validity scores in comparable settings, Ref. [19] has conditioned the generation of CIF on the space group symbols in a language model setting. Ref. [16] considered space group conditioned crystal generation using a fine-tuned generative model with space group labels. Neither approach provides exact constraints on the space group, which could yield problematic structures for large systems and highly symmetric space groups. Ref. [23] considered generating symmetric crystals in their Wyckoff representation. However, the model does not consider fractional coordinates and lattice parameters, so it requires a subsequent computational search to completely determine the crystal structure.
Figure S1(a) shows the validity of generated samples as a function of sampling temperature. One clearly sees that reduced temperature increases the validity of samples at the cost of reducing the diversity [20]. Figure S1(b) shows the novelty and uniqueness evaluated on generated samples with temperature. Novelty quantifies the proportion of new structures in the generated samples that were unseen in the training dataset. Uniqueness represents the percentage of distinct, non-redundant structures among the generated samples [123]. One sees that across different temperatures the uniqueness remains high, indicating the model does not collapse to a mode that produces duplicated samples. On the other hand, the model produces close to novel material as temperature increases, which nicely demonstrates modal covering behavior of the maximum likelihood estimation training [124]. Having a model distribution broader than the span of the dataset is crucial for material discovery.
Appendix C Discovered crystal samples with symmetric structure initialization
Table LABEL:table:samples and Table LABEL:table:pyxtal_samples list novel samples with eV/atom with CrystalFormer and PyXtal initializations respectively. Among them, Pd3Pt, MgAgPd2, TaNbRu2, Gd2HgAu, and TbMgPd2 are both discovered by CrystalFormer and PyXtal initialization. These compounds share identical crystal structures in the two tables, with lattice constants differing by less than 0.02 .